# Learning Language-guided Adaptive Hyper-modality Representation for Multimodal Sentiment Analysis

### v1

@@ Appendix A Effects of different Query, Key, and Value Settings in Fusion Transformer
$H_{hyper}^{3}$ $H_{l}^{3}$ 48.10 0.707 44.64 0.410
$H_{l}^{3}$ $H_{hyper}^{3}$ 49.42 0.683 45.73 0.404
Appendix B Analysis on Model Complexity
@@
Table 7: Analysis on model complexity. Note: the parameter of other Transformer-based methods was calculated by authors from open source code with default hyper-parameters on MOSI
@@
Appendix C Visualization of Robustness of AHL
@@
Figure 7: Visualization of the attention weights between language and visual modalities learned by the AHL for a randomly selected sample with and without a random noise on the CH-SIMS dataset. (a) The attention weights without a random noise; (b) the attention weights with a random noise. Note: darker colors indicate higher attention weights for learning
Appendix D Hyper-parameters
@@ D.1 Overview
D.2 Effects of Length Settings of Modality Feature
@@
D.3 Effects of Depth Settings of AHL
@@
D.4 Effects of Depth Settings of Fusion Transformer

### v2

@@ 4.5.5 Effects of Different Fusion Techniques
Concatenation 48.69 0.703 43.76 0.410
Addition 46.36 0.706 42.45 0.411
GRU 47.81 0.710 44.86 0.414
Tensor Fusion (TFN) 47.23 0.710 44.20 0.403
Low-rank Fusion (LMF) 46.65 0.715 45.08 0.408
Ours 49.42 0.683 45.73 0.404
4.5.6 Analysis on Model Complexity
@@
Table 8: Analysis on model complexity. Note: the parameter of other Transformer-based methods was calculated by authors from open source code with default hyper-parameters on MOSI.
@@
4.5.7 Visualization of Attention in AHL
Figure 4: Visualization of average attention weights from the last AHL layer on CH-SIMS dataset. (a) Average attention matrix $\alpha$ between language and audio modalities; (b) average attention matrix $\beta$ between language and visual modalities. Note: darker colors indicate higher attention weights for learning.
4.5.8 Visualization of Robustness of AHL
@@
Figure 5: Visualization of the attention weights between language and visual modalities learned by the AHL for a randomly selected sample with and without a random noise on the CH-SIMS dataset. (a) The attention weights without a random noise; (b) the attention weights with a random noise. Note: darker colors indicate higher attention weights for learning.
4.5.9 Visualization of Different Representations
Figure 6: Visualization of different representations in 3D space by using t-SNE.
4.5.10 Visualization of Convergence Performance
In Figure 7, we compared the convergence hehavior of ALMT with three state-of-the-art methods (i.e., MulT, MISA and Self-MM) on CH-SIMS. We choose the MAE curve for comparison as MAE indicates the model’s ability to predict fine-grained sentiment. Obviously, on the training set, although Self-MM converges the fastest, its MAE of convergence is larger than ALMT at the end of the epoch. On the validation set, ALMT seems more stable compared to other methods, while the curves of other methods show relatively more dramatic fluctuations. It demonstrates that the ALMT is easier to train and has a better generalization capability.
Figure 7: Visualization of convergence performance on the train and validation sets of CH-SIMS. (a) The comparison of MAE curves on the training set; (b) the comparison of MAE curves on the validation set. Note: the results of other methods reproduced by authors from open source code with default hyper-parameters.
5 Conclusion
In this paper, a novel Adaptive Language-guided Multimodal Transformer (ALMT) is proposed to better model sentiment cues for robust Multimodal Sentiment Analysis (MSA). Due to effectively suppressing the adverse effects of redundant information in visual and audio modalities, the proposed method achieved highly improved performance on several popular datasets. We further present rich in-depth studies investigating the reasons behind the effectiveness, which may potentially advise other researchers to better handle MSA-related tasks.
Limitations
Appendix A Hyper-parameters
@@ A.1 Overview
A.2 Effects of Length Settings of Modality Feature
@@
A.3 Effects of Depth Settings of AHL
@@
A.4 Effects of Depth Settings of Fusion Transformer